# APLe: Token-Wise Adaptive for Multi-Modal Prompt Learning

### 12 Jan 2024 v1

@@ Abstract
Pre-trained Vision-Language (V-L) models set the benchmark for generalization to downstream tasks among the noteworthy contenders. Many characteristics of the V-L model have been explored in existing research including the challenge of the sensitivity to text input and the tuning process across multi-modal prompts. With the advanced utilization of the V-L model like CLIP, recent approaches deploy learnable prompts instead of hand-craft prompts to boost the generalization performance and address the aforementioned challenges. Inspired by layer-wise training, which is wildly used in image fusion, we note that using a sequential training process to adapt different modalities branches of CLIP efficiently facilitates the improvement of generalization. In the context of addressing the multi-modal prompting challenge, we propose Token-wise Adaptive for Multi-modal Prompt Learning (APLe) for tuning both modalities prompts, vision and language, as tokens in a sequential manner. APLe addresses the challenges in V-L models to promote prompt learning across both modalities, which indicates a competitive generalization performance in line with the state-of-the-art. Preeminently, APLe shows robustness and favourable performance in prompt-length experiments with an absolute advantage in adopting the V-L models.
@@ 1 Introduction
Vision-language (V-L) pretraining is an emerging paradigm that performs image recognition tasks, such as image-recognition recognition1 (9, 19, 41), object detection object-detection1 (8, 26, 40, 35), and segmentation segmentation1 (6, 22, 31), inspirational elevating the performance in generalization ability. A novel model, CLIP CLIP (30) indicates the success of the paradigm by leveraging contrastive learning to align large-scale text-image pairs. This model introduces a simple inference process that pairs text embedding from a hand-crafted query and image embedding by calculating cosine similarity to predict the class of the image, which demonstrates the excelling generalization capability.
@@
CLIP triggers the generalization toward a new era with potential. An earlier study, CoOp COOP (44), introduced a learnable continuous prompt approach as a text input instead of relying on hand-crafted prompts shown in Fig. 1. CoOp further reveals the potential and challenge of the vision-language pretraining paradigm by delivering exceptional results on the accuracy of image classification. It also tackles the challenges of the extensive time demanded for training. Recent-coming research, MaPLe multi-modal (17), addresses the insufficient generalization performance for the downstream task if only the language prompt is explored. Thus, MaPLe exploits the prompting approach in both modalities’ prompts demonstrated in Fig. 1. The prompts are learned from both vision and language sources, while MaPLe initiates vision prompts from language counterparts by the unidirectional parsing function.
@@
Despite the effectiveness of MaPLe toward sufficiently exploiting both behaviours of both modalities to contrastive learning. Its parsing mechanism makes it infeasible to conduct prompt learning in two modalities equally due to the disparate sophistication and conditions. Such mechanisms potentially introduce challenging problems resulting in generalization capability, such as overfitting caused by the prompt length. To this end, inspired by layer-wise training and image recognition that contributes to separately extracting and processing the features from different images. This paper proposes Token-wise Adaptive for Multi-modal Prompt Learning (APLe) to adopt multi-modal prompt learning independently and sequentially to tackle the challenges and boost generation capability with the framework shown in Fig. 1.
@@
In the fundamental generalization evaluation experiments, APLe demonstrated a comparable performance with the existing methods, particularly in datasets with domain shifts. Distinctively, APLe exploits the knowledge from both modalities and illustrates the advantage of robustness by sequential prompt learning and image adaptation, while the existing approaches are insufficient or omit the image conditions and multi-modal prompt conditions. Thanks to its effectiveness, APLe attains competitive generalization capability over the baselines owing to the effective addressing of the challenges caused by image complexity and prompt length. To sum up, the contributions of the paper are summarized as follows:
@@
  • This approach conducts prompt learning with zero-shot knowledge followed by a token adaptation function to alleviate knowledge conflicts and encourage collaboration between modalities. The adaptation function encourages APLe to align the V-L representation.
Figure 2: Overview of APLe (Token-wise Adaptive for Multi-modal Prompt Learning) framework for prompting. The prompts have been trained sequentially as tokens with CLIP zero-shot knowledge and adaptation to alleviate the knowledge conflicts and prompt the synergy between modalities. APLe also deploys an image adapter to mitigate image noise and enhance image features effectively. Specifically, CE denotes the cross-entropy loss, KL denotes the Kullback-Leibler loss and $\lambda$ presents the hyper-parameter used to combine the loss functions.
@@ 2 Related Work
Vision Language Models: The multi-modalities research in exploring vision-language interaction is a core topic. At the early stage, a few attention-based approaches are introduced including BAN BAN (18), Intra-Inter Intra-Inter (10) and MCAN MCAN (39), which make contributions to the downstream tasks. While the following approaches are proposed with BERT-like architecture, ViLBERT ViLBERT (25), LXMERT LXMERT (34) and UNITER UNITER (3), expend the boundary in reasoning between the modalities. In the near past, the two prominent models, CLIP CLIP (30) and ALIGN ALIGN (14) have dominated the visual-language tasks through disparate representation learning with an impressive performance.
@@
To adopt the V-L model in downstream tasks, the conventional approaches do not adapt the current preference, which constrains the zero-shot or few-shot capability by fine-tuning and linear probing. Thanks to the supervised training with abundant data from the internet, $\sim$400M image-text pairs used in CLIP and $\sim$1B used in ALIGN, both models demonstrated the power of visual-language pre-training paradigm in zero-shot learning and few-shot learning. Inspired by CLIP and ALIGN, many following approaches achieve dominant generalization performance on specific tasks by adopting the V-L paradigm including few-shot image-recognition recognition1 (9, 19, 41), object detection object-detection1 (8, 26, 40) and segmentation segmentation1 (6, 22, 31).
@@
Prompt Design: The prompt design inspires research interest by the brand new paradigm, “pre-train, prompt, and predict”, which was used in NLP tasks NLP1 (21, 23, 24) at the early stage. The recently proposed approaches introduced the prompt learning paradigm to adopt the V-L model to downstream tasks which takes a query, as known as prompt, into the language branch to boost the generalization capability in understanding tasks. Therefore, by the effectiveness of prompt design, a rising trend of research lies on prompt design or prompt learning to further extend the potential of the V-L models and address the newly arisen challenges caused by prompt learning such as the language-only prompting COOP (44, 43, 45), vision-only prompting vision1 (15, 36, 37) and a recently proposed method, MaPLe, in multi-modal prompting multi-modal (17).
@@
Prompt Learning in Vision and Language: The early work CoOp COOP (44) exploits a learnable token into a language branch for fine-tuning CLIP. Its subsequent work Co-CoOpCOCOOP (43) further extends the boundary of generalization by exploiting conditional prompt. For the works applying prompt learning in the vision branch, Visual1 (1) is the first work to perform vision prompt in CLIP while video18 (16) introduces the prompt learning paradigm into video understanding tasks. The recent work, MaPLe multi-modal (17), proposed a framework that applies learnable tokens into both vision and language branches by coupling function. MaPLe shows an advanced performance in downstream tasks owing to the coupling function improving the alignment between vision and language representations.
@@ 3 Method
The proposed approach tackles the V-L pretraining model, CLIP, in downstream tasks. The holistic framework of our approach, APLe, is demonstrated in Fig 2. We argue that the two modalities, vision and language, equally contribute to downstream tasks while also delving into the challenges of disparate sophistication in prompt learning. To efficiently address the challenges,APLe proposes a framework that encourages prompts learned independently and sequentially from two modalities. Explicitly, APLe first attaches a Gaussian Filter image adapter to alleviate the sophistication of the image features before prompt learning, which empowers the model to effectively mitigate image noise and enhance image features across various scales. Then, APLe is structured into two blocks. A sequential token-wise knowledge training block first explores learnable prompts as tokens for training with the CLIP zero-shot model, which obtains generalization knowledge and encourages multi-modality tokens to learn knowledge independently and sequentially without interference from the other. Ultimately, the multi-modal token adaptation block promotes the synergy between the tokens.
@@ 3.1 Revisiting CLIP
CLIP presents a basic architecture to adopt the vision-language (V-L) model and addresses image recognition and classification tasks. Many existing approaches follow the architecture of CLIP COOP (44, 41). Many components used in CLIP including the neural network backbone, transformer and Vit (vision transformer), are inherited in the recent research and our approach, APLe. To adapt feature representation in CLIP, the text encoder transformed a hand-crafted query $Q$ as prompt into feature representations, and the image encoder progressed likewise to the image as features $I\in\mathbb{R}^{3\times H\times W}$ for cosine similarity calculation and generating the prediction result. The process pipeline for CLIP is demonstrated below.
Text Feature Block: Initially, the text query $Q$ is tokenized and transits into word embedding. Then, the transformer layers $\mathcal{B}$ compute word embedding and forward its result to the projection function for the cosine similarity calculation.
@@
Image Feature Block: Similar to text feature processing, the procedure excludes tokenization but appends a concatenation function for combining the class identification token $c$.
@@
Overall: The text query in the template “a photo of {class}” is transferred as $Z_{i}$, the feature $f$ of the target image and the class to predict denoted as $i\in$1,2,3…C}, the final step for cosine similarity calculation and predictions shown below:
@@ 3.2 APLe: Token-wise knowledge training and Multi-modal Token Adaptation
This paper argues that the existing methods for uni- or multi-modalities carry out the learnable prompt training with weak considering the potential and challenge of the differences in vision and language modalities. To this end, we introduce token-wise training and adaptation framework to multi-modalities prompt learning which addresses the challenges by independently and sequentially training for both prompts and optimizing the utilization of the potential of the V-L model. Image Fusion with Adapter: We argue that disparate sophistication in images caused difficulty in prompt learning. To this end, a simple and commonly seen function to obtain a lucid and authentic image feature from the original image. Inspired by the existing work in image recognition tasks (e.g. CDDFuse CDD (42)), the image $I$ in APLe is processed by Fast Fourier Transform $\mathcal{F}(\cdot)$ followed by the Gaussian filter $\mathcal{G}(\cdot)$ for suppressing image noise and enhancing subtle image features. the simple processing pipeline is shown below.
@@
Then, the image fusion operation is applied to the adjusted image and the original image where $\alpha$ is set to 0.9 The result in image fusion $\bar{I}$ is sequentially taken by Eqs.15-17 to produce the image final feature for contrastive learning. The fundamental aim of image fusion is to stabilize the variance of image features by suppressing interference rather than eliminating it. While image fusion alleviates the complexity of the image feature, the challenge is further efficiently compensated by vision tokens training in the APLe token-wise knowledge training and adaptation framework, which allows APLe to enhance its generalization performance further.
@@
By introducing fusion image $\bar{I}$ into the image feature embedding pipeline, the new image feature $\bar{f}$ is taken into Eq. 11 for contrastive learning with text feature $Z$ for predicting $p_{i}$.
Feature Embedding: In contrast to the existing approach, APLe initiates learnable prompt $D$ and $G$ for text and image respectively in a decoupling manner, which is embedded into their modalities feature by concatenation operation as seen in the Eq. 13 and Eq. 16. As the effectiveness of deep prompting, the component is maintained to APLe, which is deployed to process the newly concatenated prompts. In other settings, the length $m$ for both learnable tokens is set to 2 and the number of deep transformer layers is set with extra depth $S$. The feature embedding process is shown in the following section.
Feature Embedding in Text Feature: The encoding to text is similar to CLIP, but a learnable token $D$ is embedded.
@@
Feature Embedding in Image Feature: The image encoding corresponds to CLIP, but the image is first processed by the image adapter followed by the embedding process. Then, the result is concatenated with the learnable token $G$.
@@
In contrast with other approaches, the decoupling structure enables both vision and language prompts to learn knowledge separately and alleviate the constraint from the other, which enhances the capability of multi-modal prompts learning.
Training and Adaptation: Preeminently, APLe explicitly and effectively boost the generalization capability by conducting a distinctive and sequential training and adaptation process for prompt learning across the multiple modalities.
Sequential Token-Wise Knowledge Training: APLe train each modality sequentially and separately as embedding tokens at a time while keeping another one frozen. During the training, APLe incorporates CLIP zero-shot knowledge in the prompt learning to mitigate the potential of overfitting caused by prompt setting and escalates generalization capability using diverse combination parameters to vision and language modalities, $\lambda_{g}$ and $\lambda_{d}$ respectively, as shown in Eqs. 1819.
@@
where $\mathcal{L}_{ce}$ and $\mathcal{L}_{kl}$ are the cross-entropy loss for classification prediction and Kullback-Leibler divergence loss for quantifying the prediction disparity between the trained model and CLIP zero-shot model, which are denoted as $p_{i}$ and $p_{i}^{zs}$ respectively. While separately training the multi-modal prompt, the hyper-parameter $\lambda$ is set for combining with KL loss, while $\lambda_{g}$ and $\lambda_{d}$ are set to 0.3 for the vision prompt training and 0.5 for the language prompt training respectively.
Multi-modal Token Adaptation: Owing to the token-wise knowledge training, APLe explores the potential of both modalities in prompt learning. For further adapting the synergy of the tokens from both branches and attaining advanced generalization capability in downstream tasks, We introduce a straightforward and transparent approach that retrains the model to promote synergy between the tokens without zero-shot knowledge as the end of the pipeline of APLe.

### 23 Jan 2024 v2

@@ Abstract
Pre-trained vision-language (V-L) models set the benchmark for generalization to downstream tasks among the noteworthy contenders. Many characteristics of the V-L model have been explored in existing research including the challenge of the sensitivity to text input and the tuning process across multi-modal prompts. With the advanced utilization of the V-L model like CLIP, recent approaches deploy learnable prompts instead of hand-craft prompts to boost the generalization performance and address the aforementioned challenges. Inspired by layer-wise training, which is wildly used in image fusion, we note that using a sequential training process to adapt different modalities branches of CLIP efficiently facilitates the improvement of generalization. In the context of addressing the multi-modal prompting challenge, we propose token-wise Adaptive for multi-modal Prompt Learning (APLe) for tuning both modalities prompts, vision and language, as tokens in a sequential manner. APLe addresses the challenges in V-L models to promote prompt learning across both modalities with adaptation, which indicates a competitive generalization performance in line with state-of-the-art baselines. Preeminently, APLe shows robustness results in prompt length and overfitting experiments with an advantage by adopting CLIP zero-shot knowledge to token-wise knowledge training framework.
@@ 1 Introduction
Vision-language (V-L) pretraining is an emerging paradigm that performs computer vision tasks, such as image recognition  recognition1 (8, 18, 41), object detection object-detection1 (7, 25, 40, 37), and segmentation segmentation1 (5, 21, 30), while it inspired general application in other research 9923419 (33, 32, 34), inspirational elevating the performance in generalization ability. A novel model, CLIP CLIP (29) indicates the success of the paradigm by leveraging contrastive learning to align large-scale text-image pairs. This model introduces a simple inference process that pairs text embedding from a hand-crafted query and image embedding by calculating cosine similarity to predict the class of the image, which demonstrates the excelling generalization capability in the tasks.
@@
CLIP triggers the generalization toward a new era with potential. An earlier study, CoOp COOP (43), introduced a learnable continuous prompt approach as a text input instead of hand-crafted prompts shown in Fig. 1. CoOp further reveals the potential and challenge of the vision-language pretraining paradigm by delivering exceptional results on the accuracy of image classification, while it also tackles the challenges of the extensive time demanded for training. Recent-coming research, MaPLe multi-modal (16), argues that it is the insufficient generalization performance for the downstream task if only the language prompt is explored. Thus, MaPLe exploits the prompting learning in both modalities shown in Fig. 1, while MaPLe initiates vision prompts by transforming language prompts in a coupling function.
@@
Despite the effectiveness of MaPLe toward sufficiently exploiting both behaviors of both modalities to contrastive learning. Its coupling mechanism makes it infeasible to conduct prompt learning in two modalities equally due to the disparate sophistication and conditions. Such mechanisms potentially introduce challenging problems resulting in generalization capability, such as overfitting caused by the prompt length. To this end, inspired by layer-wise training and image fusion that contributes to separately extracting and processing the different features from images. This paper proposes token-wise Adaptive for multi-modal Prompt Learning (APLe) to adopt multi-modal prompt learning independently and sequentially to tackle the challenges and boost generation capability with the framework demonstrated in Fig. 1.
@@
In the fundamental experiments for generalization evaluation, APLe demonstrated a comparable performance with the existing methods, particularly in datasets with domain shifts. Distinctively, APLe exploits the knowledge from both modalities and illustrates the advantage of robustness by sequential prompt learning and image adaptation, while the existing approaches are insufficient or omit the image conditions and multi-modal prompt conditions. Thanks to its effectiveness, APLe attains competitive generalization capability over the baselines by effectively addressing the challenges caused by image complexity and prompt length. To sum up, the main contributions of the paper are summarized as follows:
@@
  • This approach conducts prompt learning with zero-shot knowledge followed by a token adaptation function to alleviate knowledge conflicts and encourage collaboration between modalities, thereby the adaptation function encourages APLe to align the V-L representation.
Figure 2: Overview of APLe (Token-Wise Adaptive for Multi-Modal Prompt Learning) framework for prompting. APLe first deploys an image adapter to mitigate image noise and enhance image features effectively. Then, the prompts were trained sequentially as tokens with CLIP zero-shot knowledge and adaptation to alleviate the knowledge conflicts and prompt the synergy between modalities. Specifically, CE denotes the cross-entropy loss, KL denotes the Kullback-Leibler loss and $\lambda$ presents the hyper-parameter used to combine the loss functions.
@@ 2 Related Work
Vision Language Models. The multi-modalities research in exploring vision-language interaction is a core topic. At the early stage, a few attention-based approaches are introduced including BAN BAN (17), Intra-Inter Intra-Inter (9), and MCAN MCAN (39), which make contributions to the downstream tasks. While the following approaches are proposed with BERT-like architecture, ViLBERT ViLBERT (24), LXMERT LXMERT (36) and UNITER UNITER (2), expand the boundary in reasoning between the modalities. In the near past, the two prominent models, CLIP and ALIGN ALIGN (13) have dominated the vision-language tasks through representation learning with an impressive generalization performance.
@@
To adopt the V-L model in downstream tasks, the conventional approaches do not satisfy the current preference, which constrains the zero-shot or few-shot capability by fine-tuning and linear probing. Thanks to the supervised training with abundant data from the internet, $\sim$400M image-text pairs used in CLIP and $\sim$1B used in ALIGN, both models demonstrated the power of visual-language pre-training paradigm in zero-shot learning and few-shot learning. Inspired by CLIP and ALIGN, many recent proposed approaches achieve dominant generalization performance on specific tasks by adopting the V-L paradigm into downstream tasks including few-shot image recognition, object detection, and segmentation.
@@
Prompt Design. The prompt design inspires research interest by the brand new paradigm, “pre-train, prompt, and predict”, which was used in NLP tasks NLP1 (20, 22, 23) at the early stage. The recently proposed approaches introduced the prompt learning paradigm to adopt the V-L model to downstream tasks which takes a query, as known as a prompt, into the language branch to boost the generalization capability in downstream tasks. Therefore, regarding the effectiveness of prompt design, a rising trend of research lies on prompt design or prompt learning to further extend the potential of the V-L models and address the newly arisen challenges caused by prompt learning, such as the language-only prompting model CoOp, vision-only prompting model VPT  vision1 (14) and a recently proposed method, MaPLe, in multi-modal prompting.
@@
Prompt Learning in Vision and Language. The early work CoOp exploits a learnable token into a language branch for fine-tuning CLIP. Its subsequent work Co-CoOp COCOOP (42) further extends the boundary of generalization by exploiting conditional prompt. For the works applying prompt learning in the vision branch, VPT is the first work to perform vision prompt in CLIP while video18 (15) introduces the prompt learning paradigm into video understanding tasks. The recent work, MaPLe, proposed a framework that applies learnable tokens into both vision and language branches by coupling function. MaPLe shows an advanced generalization performance owing to the coupling function improving the alignment between vision and language representations.
@@ 3 Method
The proposed approach tackles the V-L pretraining model, CLIP, in downstream tasks. We argue that the two modalities, vision and language, equally contribute to downstream tasks while also delving into the challenges of disparate sophistication in prompt learning. To efficiently address the challenges, APLe proposes a framework that encourages prompts learned independently and sequentially from two modalities. Explicitly, APLe first attaches a Gaussian filter image adapter to alleviate the sophistication of the image features before prompt learning, which empowers the model to effectively mitigate image noise and enhance image features across various scales. Then, APLe is structured into two blocks. A sequential token-wise knowledge training block first explores learnable prompts as tokens for training with the CLIP zero-shot model, which obtains generalization knowledge and encourages multi-modality tokens to learn knowledge independently and sequentially without interference from the other. Ultimately, the multi-modal token adaptation block promotes the synergy between the tokens to adopt the V-L model.
@@ 3.1 Revisiting CLIP
CLIP presents a basic architecture to adopt the vision-language model and addresses image recognition and classification tasks. Many existing approaches follow the architecture of CLIP. Many components used in CLIP including the neural network backbone, transformer and Vit (vision transformer), are inherited in the recent research and our approach, APLe. To adapt feature representation in CLIP, the text encoder transformed a hand-crafted query $Q$ with length $T$ as prompt into feature representations, and the image encoder progressed likewise to the image as features $I\in\mathbb{R}^{3\times H\times W}$ for calculating cosine similarity and generating prediction.
Text Feature Block. Initially, the text query $Q$, “a photo of {class}”, is tokenized and embedded. Then, the transformer layers $\mathcal{B}$ compute word embedding and forward its result to the projection function for the cosine similarity calculation.
@@
Image Feature Block. Similar to text feature processing, the procedure excludes tokenization but appends a concatenation function for combining the class identification token $c$.
@@
Overall. The text feature from prompts is transferred as $Z_{i}$, the image feature $f$ of the target image and the class to predict denoted as $i\in$1,2…C}, the final step of CLIP in cosine similarity calculation and predictions is shown below:
@@ 3.2 APLe: Token-Wise Knowledge Training and Multi-Modal Token Adaptation
This paper argues that the existing methods for uni- or multi-modalities carry out the learnable prompt training with weak considering the potential and challenge of the differences in vision and language modalities. To this end, we introduce token-wise training and adaptation framework to multi-modalities prompt learning which addresses the challenges by independently and sequentially training for both prompts and optimizing the utilization of the potential of the V-L model. While the holistic framework of APLe is shown in Fig. 2.
Image Fusion with Adapter. We note that disparate sophistication in images caused difficulty in prompt learning. Inspired by the existing work in image recognition tasks, a simple function is introduced to obtain a lucid and authentic image feature from the original image. The image $I$ in APLe is processed by Fast Fourier Transform $\mathcal{F}(\cdot)$ followed by the Gaussian filter $\mathcal{G}(\cdot)$ for suppressing image noise and enhancing subtle image features as the pipeline shown below:
@@
Then, the image fusion operation is applied to the adjusted image and the original image where $\alpha$ is set to 0.9. The result in image fusion $\bar{I}$ is sequentially taken by image embedding. The fundamental aim of image fusion is to stabilize the variance of image features by suppressing interference rather than eliminating it. While image fusion alleviates the complexity of the image feature, the challenge is further efficiently compensated by vision tokens training in the APLe token-wise knowledge training and adaptation framework, which allows APLe to enhance its generalization performance further.
@@
By introducing fusion image $\bar{I}$ into the image feature embedding pipeline, the new image feature $\bar{f}$ is taken into Eq. 11 for contrastive learning and predicting $p_{i}$ with text feature $Z$.
Feature Embedding. In contrast to the existing approach, APLe initiates learnable prompt $D$ and $G$ for text and image respectively, in a decoupling manner, which is embedded into their modalities feature by concatenation operation as seen in the Eq. 13 and Eq. 16. As the effectiveness of deep prompting, the component is maintained to APLe, which is deployed to process the newly concatenated prompts. In other settings, the length $m$ for both learnable tokens is set to 2 and the number of deep transformer layers is set with extra depth $S=8$, while $d_{lang}=512$, $d_{vis}=768$, and $d_{vis-lang}=512$.
Feature Embedding in Text Feature. The encoding to text is similar to CLIP, but a learnable token $D$ is embedded.
@@
Feature Embedding in Image Feature. The image encoding corresponds to CLIP, but the image is first processed by the image fusion followed by the embedding process. Then, the result is concatenated with the learnable vision token $G$.
@@
In contrast with other approaches, the decoupling structure enables both vision and language prompts to learn knowledge separately, and alleviates the constraint from the other, which enhances the capability of multi-modal prompts learning.
Training and Adaptation. Preeminently, APLe explicitly and effectively boosts the generalization capability by conducting a distinctive and sequential training and adaptation process for prompt learning across multiple modalities.
Sequential Token-Wise Knowledge Training. APLe trains each modality prompt sequentially and separately as tokens at a time while keeping another one frozen. During the training, APLe incorporates CLIP zero-shot knowledge in the prompt learning to escalate generalization capability and mitigate the potential of overfitting due to prompt setting by using diverse combination parameters to language and vision prompts, $\lambda_{d}$ and $\lambda_{g}$ respectively, as demonstrated in Eqs. 1819.
@@
where $\mathcal{L}_{ce}$ and $\mathcal{L}_{kl}$ are the cross-entropy loss for classification prediction and Kullback-Leibler divergence loss for quantifying the prediction disparity between the trained model and CLIP zero-shot model, which are denoted as $p_{i}$ and $p_{i}^{zs}$ respectively. While separately training the modalities prompts, the hyper-parameter $\lambda$ is set for combining with KL loss, while $\lambda_{d}$ and $\lambda_{g}$ are 0.5 for the language prompt training and set to 0.3 for the vision prompt training respectively.
Multi-Modal Token Adaptation. Owing to the token-wise knowledge training, APLe explores the potential of both modalities in prompt learning. For further adapting the synergy of the tokens from both branches and attaining advanced generalization capability in downstream tasks, we introduce a straightforward and transparent approach that retrains the model to promote synergy between the tokens without zero-shot knowledge as the end of the pipeline of APLe.